### v1

@@ Abstract
We present the first hadron-level predictions at MEPS@NLO precision for both neutral and charged current DIS at the EIC and the first application of the MEPS@NLO technique to charged current DIS in general. We critically examine state-of-the-art inclusive predictions using multi-leg merging techniques, contrasting perturbative and non-perturbative uncertainties. We study typical kinematic DIS observables as well as jet measurements and 1-jettiness with realistic cuts implied by expected and past detector resolution. On the perturbative side, we see large corrections towards small virtualities and Bjorken-$x$, which can be captured by higher-multiplicity matrix elements and the merging procedure. Non-perturbative effects, while negligible in most jet observables, can reach similar size as the perturbative uncertainties around the peak of the 1-jettiness distributions especially at low values of $Q^{2}$.
@@ 1 Introduction
Lepton-hadron collision experiments, as conducted for example at HERA, are crucial for probing hadron structure [1, 2, 3]. These interactions occur via the exchange of at least one electroweak boson. When the virtuality of this mediator is sufficiently large, i.e., $Q^{2}\gg m_{P}^{2}\sim 1\,\text{GeV}^{2}$, the interaction enters the Deeply Inelastic Scattering (DIS) regime. In this limit, the factorization of the cross-section into matrix elements and parton distribution functions (PDFs) has been proven explicitly [4, 5], hence allowing the application of perturbative techniques. Specifically, in the case of neutral-current (NC) DIS mediated by photons and Z-bosons, in contrast to charged-current (CC) DIS mediated by W-bosons, measurements can be performed fully inclusively with respect to the hadronic final state, providing a high-precision probe of the proton’s partonic structure.
@@
In recent years, lepton-hadron scattering has garnered more attention again due to the upcoming planned for the Electron-Ion Collider (EIC) at Brookhaven National Laboratory. The EIC aims for unprecedented precision in determining the spatial and transverse-momentum distributions of partons in the proton [6, 7]. Maximizing the available statistics for experimental analyses requires lowering the virtuality cut-off, as the cross-section scales as $1/Q^{4}$. However, at lower $Q^{2}$, additional scales – such as jet transverse momenta – can become dominant, leading to increased theoretical uncertainties. This challenge is addressed by the computation of higher-order corrections in fixed-order (FO) perturbation theory, particularly via real radiation corrections in QCD. While current FO calculations extend up to N3LO [8, 9, 10], further improvements are computationally prohibitive in the foreseeable future. Additionally, the lower center-of-mass energy than, e.g. , at HERA shifts the attention toward smaller values of $Q^{2}$ as well as non-perturbative corrections which should approximately be of the order $\Lambda_{\mathrm{QCD}}/Q$ or $\Lambda^{2}_{\mathrm{QCD}}/Q^{2}$.
@@
An alternative approach is multi-jet merging, which combines tree-level matrix elements of varying final-state multiplicities and parton shower evolution. This method was initially developed at leading order (LO) [11, 12] and later extended to next-to-leading order (NLO) [13, 14]. Multi-jet merging for DIS was recently also studied at LO within the P YTHIA event generator [15], too. Although not a complete higher-order correction, this approach effectively captures the dominant effects at small virtualities, in particular in the range $Q^{2}\simeq\mathcal{O}(1\text{-}10\,\text{GeV}^{2})$.
@@
Most previous studies of inclusive DIS phenomenology for the EIC have relied on LO calculations, primarily using P YTHIA  [16, 17, 18, 19, 20, 21, 22]. Notably, Ref. [16] explored the full-$Q^{2}$ phase space, including the low-$Q^{2}$ photoproduction regime at LO. DIS has also been recently studied at matched NLO accuracy in the P OWHEG formalism in Refs. [23, 24, 25]. Several other tools [26, 27] are available to compute DIS observables at NNLO accuracy for the EIC. Furthermore, jet predictions in the photoproduction regime were computed at NLO in Ref. [28], see also Ref. [29].
@@
In this work, we present a study of precision DIS phenomenology at the EIC using the S HERPA event generator. Specifically, we employ its automated matching and merging capabilities to produce the most precise, fully exclusive hadron-level predictions to date. We analyze both traditional DIS observables and hadronic event shapes, including 1-jettiness, jet multiplicities, and leading-jet transverse momentum. To assess theoretical uncertainties, we perform independent evaluations of perturbative and non-perturbative effects, identifying limitations in theoretical predictions relevant to EIC analyses.
@@
We structure our paper as follows: In Sec. 2, we introduce our notation for DIS kinematics and define the observables to be studied. Sec. 3 details our simulation framework within S HERPA , and Sec. 4 presents a validation of the CC DIS simulation against ZEUS data. We then analyze NC and CC DIS events at the EIC in Sec. 5, before concluding in Sec. 6.
@@ 2 Overview of DIS kinematics and analyzed observables
This results in collection of hadrons with momentum $p_{X}$ plus a recoiling fermion momentum $p^{\prime}$ which can be either another electron, $e^{-}$, or a neutrino, $\nu_{e}$, if the interaction is driven by neutral or charged current, respectively***While we focus on the case of electron beams here, the described methods are of course likewise available for positron beams. In both cases the virtual boson carries momentum $q=p-p^{\prime}$. When the momentum transfer is large enough, the virtual boson acts as a probe of the internal structure of the proton, which is dominated by color-confined strong interactions. Thus, measurements of this family of scattering process are an excellent experimental handle on the behavior of strong nuclear forces and the nature of the elementary degrees of freedom of the hadronic side of the Standard Model.
It is customary to parameterize the scattering using two out of three of the following invariants
@@
Bjorken $\text{Bjorken~{}x}\;\;\;x$ $=\frac{Q^{2}}{2P\cdot q}\ ,$ (2b)
and $\text{and inelasticity}\;\;\;y$ $=\frac{P\cdot q}{P\cdot p}\ \ .$ (2c)
The observable $Q^{2}$ defines the energy scale of the boson probing the proton structure, while Eq. (2b) defines the lowest fraction of momentum that a struck quark or gluon inside the proton can carry. Finally, Eq. (2c) quantifies the energy loss of the electron. This choice of variables is especially relevant in the neutral current case as it allows to fully characterize the cross-section from measurements of the final state lepton only. Indeed, one of the main goals of the EIC program is a high precision determination of the proton structure and its dependence on $Q^{2}$ [6], this motivates theoretical effort in improving the prediction for the same observables.
Beside the inclusive cross-section, we can learn more about the features of the strong interactions, e.g. the value of strong coupling $\alpha_{\text{s}}$ or the details of the QCD radiation patterns, by studying directly properties of the ensemble of hadrons emerging in the final state. Event shapes observables are one way to approach this task as they evaluate the geometrical distribution of particles in the final state. Specifically for this work, we consider the 1-jettiness distribution,
@@
[…]
Finally, given a definition of jets, one more handle on the behavior of QCD radiation is offered by the internal structure of the jet themselves. Analyses of jet substructure are a tool extensively used at the LHC [35, 36, 37, 38]. A common goal is to remove soft radiation at wide angles from the jet axis, which is typically dominated by contamination from sources outside the jet and hadronization corrections. This can be achieved by the so called soft-drop grooming technique [39]. While originally introduced in the context of jets at hadron colliders, and the LHC specifically, it has since been explored in many other contexts such as event shapes at hadron [40] as well as lepton colliders [41, 42], in Higgs decays [43] and for jets measured at lower energy experiments such as RHIC [44]. Here, we consider the application in DIS described in [45] to soft-drop on top of the C ENTAURO clustering (see above). Related observables have recently been measured in DIS for the first time [46]. Briefly, this entails first clustering the hadronic ensemble $X$ with the standard algorithm and then step backwards through the jet clustering history. At each branching, if
@@
where $z_{\text{cut}}$ defines the cut-off parameter for the procedure, the softer subjet is discarded and the procedure continues along the harder branch. The algorithm terminates if either Eq. (5) is not met or there is only one particle left in the jet. After this grooming procedure is completed the surviving particles can be used to compute any hadronic observable. For this work, we will consider the groomed 1-jettiness:
@@ 3 Event Generation for DIS in S HERPA
For our hadron-level predictions, we use the S HERPA v3 [47] event generator. It provides matrix elements at tree level through the internal generators A MEGIC [48] and C OMIX [49]. Virtual corrections for the DIS process are likewise included internally. We use S HERPA ’s default css[50] based on Catani–Seymour dipoles [51]. It can be matched to NLO matrix elements using the MC@NLO [52] method. Matrix elements with different multiplicities are combined using the CKKW merging method [53]. Events can be hadronized using the cluster hadronization model [54] as implemented in S HERPA [55, 56] or alternatively the Lund model using an interface to P YTHIA 8 [57]. The events are analyzed using S HERPA ’s interface to R IVET  [58].
@@ 3.1 MEPS@LO, MC@NLO and MEPS@NLO
The automation of matching and merging for DIS in S HERPA has been presented in [12]. Matching to higher order matrix elements is achieved through S HERPA ’s implementation of the MC@NLO prescription [59]. Merging in the CKKW method is achieved by applying a jet criterion at the matrix-element level and restricting the available phase space for the parton shower. To maintain the logarithmic accuracy, a backwards clustering has to be performed to dress matrix elements with appropriate Sudakov factors and evaluate $\alpha_{\text{s}}$ at the correct scale. The techniques used in this work are known as MEPS@LO [11] and MEPS@NLO [13, 14], and we refer to the respective publications for the details of the procedures and subtleties related to the use of NLO matrix elements.
In our calculations, we compute single- and di-jet production at NLO and three- and four-jet production at LO, i.e.
@@
for CC DIS at MEPS@NLO, and analogous simplified processes for MEPS@LO, MC@NLO and LO predictions. In the neutral-current interactions, we treat the light quarks as massless and additionally compute and merge in processes with massive $c$- and $b$-quarks at LO. In the parton shower, $c$ and $b$ quarks are likewise treated using massive splitting functions[60]. For the charged-current simulation, we work with 4 massless quarks. The $b$ does not contribute to jet production since we assume a diagonal CKM matrix, and is anyway heavily suppressed by the PDF.
@@
Figure 1: Sketch of an example Feynman diagram with three final-state partons in DIS. Depending on the hierarchy of the internal momenta, the event is clustered back to an underlying $2\to 2$ core process. Only if $q^{2}\gg k_{1}^{2},\,k_{2}^{2}$ holds, is the core process DIS-like.
@@
    virtual photon exchange, i.e. $ej\to ej$, where $\mu_{\mathrm{core}}^{2}=Q^{2}$,
@@
    interaction of the virtual photon with a QCD parton, i.e. $\gamma^{*}j\to j_{1}j_{2}$, with $\mu^{2}_{\mathrm{core}}=m_{\perp,1}m_{\perp,2}$ defined as the product of the two jet transverse masses $m_{\perp,i}=\sqrt{m^{2}_{i}+p_{\perp,i}^{2}}$ relative to the beam axis,
@@
    pure QCD channels, i.e. $jj\to jj$, where $\mu^{2}_{\mathrm{core}}=-\frac{1}{\sqrt{2}}\left(s^{-1}+t^{-1}+u^{-1}\right)^{-1}$ is a scaled harmonic mean of the Mandelstam variables $s,t,u$.
The factorization and shower starting scales are then set to the core scale, i.e. $\mu_{\mathrm{F}}=\mu_{\mathrm{R}}=\mu_{\mathrm{Q}}=\mu_{\mathrm{core}}$. For higher multiplicities, the renormalization scale is determined according to the clustering algorithm [11] with $\mu_{\mathrm{core}}$ as scale of the core $2\to 2$ process.
The merging scale $Q_{\mathrm{cut}}$, i.e. the separation of the phase space into a region filled by additional legs in the matrix elements instead of the parton shower, is dynamically given by
@@
using $\bar{Q}_{\mathrm{cut}}=25\,\text{GeV}$ and $S_{\mathrm{DIS}}=0.4$. The former avoids unnecessary multi-leg matrix elements, while the latter ensures a decrease of the merging scale towards smaller virtualities, therefore improving the description in this region.
For the parton densities, we use the NNPDF30_nlo_as_0118 set [61] using L HA P DF  [62]. L HA P DF is also used for the $\alpha_{\text{s}}$ evolution, ensuring consistency with the PDF set and in particular implying $\alpha_{\text{s}}(M_{Z})=0.118$.

### v2

@@ Abstract
We present state-of-the-art hadron-level predictions for the deep-inelastic scattering process at next-to-leading-order precision for several multiplicities, consistently merged in one sample. For the first time at this level of accuracy, we consider both neutral and charged current deep-inelastic scattering at the Electron-Ion Collider, and present the first application of consistent next-to-leading-order merging to charged current deep-inelastic scattering in general. We critically examine inclusive predictions using multileg merging techniques, contrasting perturbative and nonperturbative uncertainties. Further, we study typical kinematic deep-inelastic scattering observables as well as jet measurements and 1-jettiness with realistic cuts implied by expected and past detector resolution. On the perturbative side, we see large corrections toward small virtualities and Bjorken-$x$, which can be captured by higher-multiplicity matrix elements and the merging procedure. Nonperturbative effects, while negligible in most jet observables, can reach similar size as the perturbative uncertainties around the peak of the 1-jettiness distributions especially at low values of $Q^{2}$.
@@ 1 Introduction
Lepton-hadron collision experiments, as conducted for example at HERA, are crucial for probing hadron structure [1, 2, 3]. These interactions occur via the exchange of at least one electroweak boson. When the virtuality of this mediator is sufficiently large, i.e., $Q^{2}\gg m_{P}^{2}\sim 1\,\text{GeV}^{2}$, the interaction enters the deeply inelastic Scattering (DIS) regime. In this limit, the factorization of the cross section into matrix elements and parton distribution functions (PDFs) has been proven explicitly [4, 5], hence allowing the application of perturbative techniques. Specifically, in the case of neutral-current (NC) DIS mediated by photons and Z-bosons, in contrast to charged-current (CC) DIS mediated by W-bosons, measurements can be performed fully inclusively with respect to the hadronic final state, providing a high-precision probe of the proton’s partonic structure.
@@
In recent years, lepton-hadron scattering has garnered more attention again due to the upcoming planned for the Electron-Ion Collider (EIC) at Brookhaven National Laboratory. The EIC aims for unprecedented precision in determining the spatial and transverse-momentum distributions of partons in the proton [6, 7]. Maximizing the available statistics for experimental analyses requires lowering the virtuality cutoff, as the cross section scales as $1/Q^{4}$. However, at lower $Q^{2}$, additional scales – such as jet transverse momenta – can become dominant, leading to increased theoretical uncertainties. This challenge is addressed by the computation of higher-order corrections in fixed-order (FO) perturbation theory, particularly via real radiation corrections in QCD. While current FO calculations extend up to next-to-next-to-next-to-leading order (N3LO) [8, 9, 10], further improvements are computationally prohibitive in the foreseeable future. Additionally, the lower center-of-mass energy than, e.g. , at HERA shifts the attention toward smaller values of $Q^{2}$ as well as nonperturbative corrections which should approximately be of the order $\Lambda_{\mathrm{QCD}}/Q$ or $\Lambda^{2}_{\mathrm{QCD}}/Q^{2}$.
@@
An alternative approach is multijet merging, which combines tree-level matrix elements of varying final-state multiplicities and parton shower evolution. This method was initially developed at leading order (LO) [11, 12] and later extended to next-to-leading-order (NLO) [13, 14]. Multijet merging for DIS was recently also studied at LO within the P YTHIA event generator [15], too. Although not a complete higher-order correction, this approach effectively captures the dominant effects at small virtualities, in particular in the range $Q^{2}\simeq\mathcal{O}(1\text{-}10\,\text{GeV}^{2})$.
@@
Most previous studies of inclusive DIS phenomenology for the EIC have relied on LO calculations interfaced to parton showers (LO+PS), primarily using P YTHIA  [16, 17, 18, 19, 20, 21, 22]. Notably, Ref. [16] explored the full-$Q^{2}$ phase space, including the low-$Q^{2}$ photoproduction regime at LO. DIS has also been recently studied at matched NLO accuracy in the P OWHEG formalism in Refs. [23, 24, 25]. Several other tools [26, 27] are available to compute DIS observables at next-to-next-to-leading-order (NNLO) accuracy for the EIC. Furthermore, jet predictions in the photoproduction regime were computed at NLO in Ref. [28]; see also Ref. [29].
LO+PS calculations will provide only a leading-order description of the inclusive DIS kinematics (i.e. $Q^{2}$ and $x_{Bj}$ spectra, see Sec. 2 for details) and describe additional jets within the parton shower approximation***Since we are mainly focused on the inclusion of higher order matrix elements here, we choose to largely avoid a full discussion of the formal accuracy achieved by the parton shower. For example the discussion in [30] (see also [31] for a DIS specific discussion) indicates that the default S HERPA dipole shower used here asymptotically resums leading logarithms (LL) of event shapes at leading color, while breaking next-to-leading logarithms due to its recoil scheme. On the other hand, parton showers include important physical constrains such as exact momentum conservation and unitarity [32]. Within the S HERPA framework we use here, a shower, dubbed A LARIC , satisfying the recoil safety requirement for NLL resummation has been implemented [33, 34] and equipped with the same merging techniques at LO [35] for initial- and final state evolution and matching and merging at NLO for final state emissions [36]. In practical terms it has shown similar performance to the shower used here for the cases where A LARIC is available so far.. NLO calculations matched to parton showers will naturally describe the inclusive kinematics at NLO accuracy, and observables related to one additional emission, such as the $p_{T}$ spectrum of an additional jet, at leading order. Observables sensitive to more emissions are again described in the parton shower approximation. Fixed-order calculations are available at higher perturbative orders; however, there are no automated matching schemes available for these so far. Without matching to a parton shower, and subsequent hadronization via a dedicated model, there are no methods to relate such a calculation to a fully differential hadronic final state. Merging NLO (for higher multiplicities LO) matrix elements, as described in this manuscript, allows us to describe observables sensitive to further emissions, for example the $p_{T}$-spectrum of the $n$th jet, at NLO (LO) accuracy, while maintaining the ability to include a parton shower and hence a fully differential distribution for arbitrary observables.
@@
In this work, we present a study of precision DIS phenomenology at the EIC using the S HERPA event generator. Specifically, we employ its automated matching and merging capabilities to produce the most precise, fully exclusive hadron-level predictions to date. We analyze both traditional DIS observables and hadronic event shapes, including 1-jettiness, jet multiplicities, and leading-jet transverse momentum. To assess theoretical uncertainties, we perform independent evaluations of perturbative and nonperturbative effects, identifying limitations in theoretical predictions relevant to EIC analyses.
@@
We structure our paper as follows. In Sec. 2, we introduce our notation for DIS kinematics and define the observables to be studied. Section 3 details our simulation framework within S HERPA , and Sec. 4 presents a validation of the CC DIS simulation against ZEUS data. We then analyze NC and CC DIS events at the EIC in Sec. 5, before concluding in Sec. 6.
@@ 2 Overview of DIS kinematics and analyzed observables
This results in collection of hadrons with momentum $p_{X}$ plus a recoiling fermion momentum $p^{\prime}$ which can be either another electron, $e^{-}$, or a neutrino, $\nu_{e}$, if the interaction is driven by neutral or charged current, respectivelyWhile we focus on the case of electron beams here, the described methods are of course likewise available for positron beams. In both cases the virtual boson carries momentum $q=p-p^{\prime}$. When the momentum transfer is large enough, the virtual boson acts as a probe of the internal structure of the proton, which is dominated by color-confined strong interactions. Thus, measurements of this family of scattering process are an excellent experimental handle on the behavior of strong nuclear forces and the nature of the elementary degrees of freedom of the hadronic side of the Standard Model.
It is customary to parametrize the scattering using two out of three of the following invariants
@@
Bjorken $x\displaystyle\text${Bjorken~x}\;\;\;x $=\frac{Q^{2}}{2P\cdot q}\ ,$ (2b)
$\text{and inelasticity}\;\;\;y$ $=\frac{P\cdot q}{P\cdot p}\ \ .$ (2c)
The observable $Q^{2}$ defines the energy scale of the boson probing the proton structure, while Eq. (2b) defines the lowest fraction of momentum that a struck quark or gluon inside the proton can carry. Finally, Eq. (2c) quantifies the energy loss of the electron. This choice of variables is especially relevant in the neutral current case as it allows to fully characterize the cross section from measurements of the final-state lepton only. Indeed, one of the main goals of the EIC program is a high-precision determination of the proton structure and its dependence on $Q^{2}$ [6], this motivates theoretical effort in improving the prediction for the same observables.
Beside the inclusive cross section, we can learn more about the features of the strong interactions, e.g., the value of strong coupling $\alpha_{\text{s}}$ or the details of the QCD radiation patterns, by studying directly properties of the ensemble of hadrons emerging in the final state. Event shapes observables are one way to approach this task as they evaluate the geometrical distribution of particles in the final state. Specifically for this work, we consider the 1-jettiness distribution,
@@
where $\Delta_{ij}^{2}=\left(y_{i}-y_{j}\right)^{2}+(\phi_{i}^{2}-\phi_{j}^{2})$ is the distance in the azimuthal plane between the two hadrons and $y$, $\phi$ and $k_{t}$ are the rapidity, azimuthal angle and transverse momentum relative to the beam axis of each hadron, respectively. All quantities here are defined in the Breit frame for our purposesThe Breit frame, or brick wall frame, is defined as the frame in which the momentum transfer is $q=Q\left(0,\,0,\,0,\,-1\right),$ this provides a natural separation hemispheres between the leptonic and hadronic side of DIS at Born level. . The parameter $R$ controls the expected average size of the jet. All these tools allow for the characterization of events in terms of leading jet transverse momentum and pseudorapidity, as well as the overall jet multiplicity and invariant masses of di-/trijet systems.
Finally, given a definition of jets, one more handle on the behavior of QCD radiation is offered by the internal structure of the jet themselves. Analyses of jet substructure are a tool extensively used at the LHC [42, 43, 44, 45]. A common goal is to remove soft radiation at wide angles from the jet axis, which is typically dominated by contamination from sources outside the jet and hadronization corrections. This can be achieved by the so called soft-drop grooming technique [46]. While originally introduced in the context of jets at hadron colliders, and the LHC specifically, it has since been explored in many other contexts such as event shapes at hadron [47] as well as lepton colliders [48, 49], in Higgs decays [50] and for jets measured at lower-energy experiments such as RHIC [51]. Here, we consider the application in DIS described in Ref. [52] to soft drop on top of the C ENTAURO clustering (see above). Related observables have recently been measured in DIS for the first time [53]. Briefly, this entails first clustering the hadronic ensemble $X$ with the standard algorithm and then stepping backwards through the jet clustering history. At each branching, if
@@
where $z_{\text{cut}}$ defines the cutoff parameter for the procedure, the softer subjet is discarded and the procedure continues along the harder branch. The algorithm terminates if either Eq. (5) is not met or there is only one particle left in the jet. After this grooming procedure is completed the surviving particles can be used to compute any hadronic observable. For this work, we will consider the groomed 1-jettiness:
@@ 3 Event generation for DIS in S HERPA
For our hadron-level predictions, we use the S HERPA [54] event generator. It provides matrix elements at tree level through the internal generators A MEGIC [55] and C OMIX [56]. Virtual corrections for the DIS process are likewise included internally. We use S HERPA ’s default css[57] based on Catani-Seymour dipoles [58]. It can be matched to NLO matrix elements using the MC@NLO [59] method. Matrix elements with different multiplicities are combined using the CKKW merging method [60]. Events can be hadronized using the cluster hadronization model [61] as implemented in S HERPA [62, 63] or alternatively the Lund model using an interface to P YTHIA 8 [64]. The events are analyzed using S HERPA ’s interface to R IVET  [65]. For the parton densities, we use the NNPDF30_nlo_as_0118 set [66] using L HA P DF  [67]. L HA P DF is also used for the $\alpha_{\text{s}}$ evolution, ensuring consistency with the PDF set and in particular implying $\alpha_{\text{s}}(M_{Z})=0.118$.
@@ 3.1 MEPS@LO, MC@NLO and MEPS@NLO
The automation of matching and merging for DIS in S HERPA has been presented in [12]. Matching to higher-order matrix elements is achieved through S HERPA ’s implementation of the MC@NLO prescription [68]. Merging in the CKKW method is achieved by applying a jet criterion at the matrix-element level and restricting the available phase space for the parton shower. To maintain the logarithmic accuracy, a backward clustering has to be performed to dress matrix elements with appropriate Sudakov factors and evaluate $\alpha_{\text{s}}$ at the correct scale. The techniques used in this work are known as MEPS@LO [11] and MEPS@NLO [13, 14], and we refer to the respective publications for the details of the procedures and subtleties related to the use of NLO matrix elements.
In our calculations, we compute single- and dijet production at NLO and three- and four-jet production at LO, i.e. ,
@@
for CC DIS at MEPS@NLO, and analogous simplified processes for MEPS@LO, MC@NLO and LO predictions. In the neutral-current interactions, we treat the light quarks as massless and additionally compute and merge in processes with massive $c$ and $b$ quarks at LO. In either case we use the five-flavor NNPDF30_nlo_as_0118 PDF set. In the parton shower, $c$ and $b$ quarks are likewise treated using massive splitting functions [69]. For the charged-current simulation, we work with four massless quarks. The $b$ does not contribute to jet production since we assume a diagonal CKM matrix, and is anyway heavily suppressed by the PDF.
@@
Figure 1: Sketch of an example Feynman diagram with three final-state partons in DIS. Depending on the hierarchy of the internal momenta, the event is clustered back to an underlying $2\to 2$ core process. Only if $q^{2}\gg k_{1}^{2},\,k_{2}^{2}$ holds is the core process DIS-like.
@@
    virtual photon exchange, i.e. , $ej\to ej$, where $\mu_{\mathrm{core}}^{2}=Q^{2}$;
@@
    interaction of the virtual photon with a QCD parton, i.e. , $\gamma^{*}j\to j_{1}j_{2}$, with $\mu^{2}_{\mathrm{core}}=m_{\perp,1}m_{\perp,2}$ defined as the product of the two jet transverse masses $m_{\perp,i}=\sqrt{m^{2}_{i}+p_{\perp,i}^{2}}$ relative to the beam axis,
@@
    pure QCD channels, i.e. , $jj\to jj$, where $\mu^{2}_{\mathrm{core}}=-\frac{1}{\sqrt{2}}\left(s^{-1}+t^{-1}+u^{-1}\right)^{-1}$ is a scaled harmonic mean of the Mandelstam variables $s,t,u$.
The factorization and shower starting scales are then set to the core scale, i.e. , $\mu_{\mathrm{F}}=\mu_{\mathrm{R}}=\mu_{\mathrm{Q}}=\mu_{\mathrm{core}}$. For higher multiplicities, the renormalization scale is determined according to the clustering algorithm [11] with $\mu_{\mathrm{core}}$ as the scale of the core $2\to 2$ process.
The merging scale $Q_{\mathrm{cut}}$, i.e. , the separation of the phase space into a region filled by additional legs in the matrix elements instead of the parton shower, is dynamically given by
@@
using $\bar{Q}_{\mathrm{cut}}=25\,\text{GeV}$ and $S_{\mathrm{DIS}}=0.4$. The former avoids unnecessary multileg matrix elements, while the latter ensures a decrease of the merging scale towards smaller virtualities, therefore improving the description in this region.